OpenStudy (anonymous):

PLZ HELP ME!!! I WILL GIVE U A MEDAL TO THE BEST ANSWER: What is thirty−six percent of fifty? one point four zero point seven two eighteen seventy−two

OpenStudy (anonymous):

Plz help me i will give u a medal

OpenStudy (anonymous):

plz

OpenStudy (kagıtucak):

\[50\times \frac{ 36 }{100}=18\]

OpenStudy (anonymous):

so it is 18?

OpenStudy (kagıtucak):

exactly

OpenStudy (anonymous):

kk thanks can u help me with one more?

OpenStudy (anonymous):

@kagıtucak help with another one?

OpenStudy (kagıtucak):

ok send it

OpenStudy (anonymous):

kk thanks

OpenStudy (anonymous):

fourteen is what percent of twenty−eight? two percent five percent twenty percent fifty percent

OpenStudy (kagıtucak):

\[28\times \frac{ x }{ 100 }=14 \]

OpenStudy (anonymous):

so what is the answer?

OpenStudy (kagıtucak):

\[x=14\times100\div28=50\]

OpenStudy (kagıtucak):

so fifty percent

OpenStudy (anonymous):

kk thansk

OpenStudy (anonymous):

and plz watch my youtube chhanel